### v1

# Lower bounds for sphere packing in arbitrary norms

Abstract.
We show that in any $d$-dimensional real normed space, unit balls can be packed with density at least
$\frac{(1-o(1))d\log d}{2^{d+1}},$
improving a result of Schmidt from 1958 by a logarithmic factor and generalizing the recent result of Campos, Jenssen, Michelen, and Sahasrabudhe in the $\ell_{2}$ norm. Our main tools are the graph-theoretic result used in the $\ell_{2}$ construction and recent progress on the Bourgain slicing problem.
1. Introduction
The question of how densely congruent spheres may be packed in $d$-dimensional space is a classical one. The trivial bound, that at least a $2^{-d}$ proportion of space may be covered, was first improved by a factor of $2$ by Minkowski [Min05] in 1905. The first asymptotic improvement (i.e. by more than a constant factor) on the trivial bound was due to Rogers [Rog47], who gave a lower bound of the form $c(1-o(1))d2^{-d}$ for $c=2/e$. This constant was improved to $c=2$ in 1992 by Ball [Bal92] and to $c=65963$ by Venkatesh in 2012 [Ven13]. Very recently, Campos, Jenssen, Michelen, and Sahasrabudhe [CJMS23] gave an asymptotic improvement, replacing the constant $c$ by $\frac{1}{2}\log d$ [^1] Here and throughout, $\log$ denotes the natural logarithm.
[…]
Our main result extends the result of [CJMS23] to arbitrary norms, with the same asymptotic improvement and the same constant factor. This improves upon the previous best-known bound by a logarithmic factor. We will phrase our results in terms of packing centrally symmetric convex sets, as our arguments are more natural in this language. [^2] Compact centrally symmetric convex sets and norms on $\mathbb{R}^{d}$ are in bijection: the unit ball $\{x\in\mathbb{R}^{d}\colon\|x\|\leq 1\}$ is centrally symmetric, compact, and convex, while if $K$ is centrally symmetric, compact, and convex, then $\|x\|=\inf\{\lambda\geq 0\colon x\in\lambda K\}$ is a norm on $\mathbb{R}^{d}$. Given a compact set $K\subset\mathbb{R}^{d}$, let $\delta_{T}(K)$ be the maximum density of a packing of translates of $K$ in $\mathbb{R}^{d}$.
Theorem 1.1.
For each compact, centrally symmetric convex set $K\subset\mathbb{R}^{d}$,
$\delta_{T}(K)\geq(1-o(1))\frac{d\log d}{2^{d+1}},$
where the $o(1)$ term tends to zero with the dimension $d$, irrespective of the particular set $K$.
We quickly mention the question of finding upper bounds. In the $\ell_{2}$ case, the best upper bound is of the shape $2^{-(0.599\ldots-o(1))d}$, due to Kabatjanskiĭ and Levenšteĭn [KL78] with constant-factor improvements by Cohn and Zhao [CZ14] and Sardari and Zargar [SZ23]. Some sets $K$ (such as the hypercube) can of course be packed much more tightly, so there is no general-purpose upper bound in the setting of Theorem 1.1. We are not aware of any centrally symmetric convex set $K\subset\mathbb{R}^{d}$ whose translational packing density is known to be strictly less than the best known upper bound for the $\ell_{2}$ unit ball.
It is also interesting to consider translational packings of shapes which are not centrally symmetric. Via a result of Rogers and Shephard [RS57], we may the following from Theorem 1.1.
Corollary 1.2.
For each compact, convex set $K\subset\mathbb{R}^{d}$,
$\delta_{T}(K)\geq(1-o(1))\frac{\sqrt{\pi}d^{3/2}\log d}{2^{2d+1}}.$
We now make some comments about our approach. Historically, most of the arguments achieving dense packings of convex shapes have constructed lattice packings, i.e. packings in which the set of centers forms a lattice in $\mathbb{R}^{d}$. In contrast, the packing given in [CJMS23] is amorphous, lacking any global structure. In the case of the $\ell_{2}$ unit ball, densities asymptotic to $cd2^{-d}$ were first achieved for amorphous packings in 2004 by Krivelevich, Litsyn, and Vardy [KLV04], with constant-factor improvements by Jenssen, Joos, and Perkins [JJP19] and Fernández, Kim, Liu, and Pikhurko [FKLP23]. For $\ell_{p}$ unit balls and some generalizations (superballs), Xie and Ge [XG22] recently extended the techniques of [JJP19] to give lower bounds of the form $cd2^{-d}$. To our knowledge, ours is the first work achieving densities asymptotically greater than $2^{-d}$ for other convex bodies using amorphous packings.
In [KLV04], the analysis of the amorphous packing is via tools from graph theory. Specifically, the authors use a result of Ajtai, Komlós, and Szemerédi [AKS80] that a graph with few triangles must have a large independent set. The improvement of Campos, Jenssen, Michelen, and Sahasrabudhe comes from a novel and more refined graph-theoretic result, which allows them to obtain an extra factor of $\log d$ in the size of the independent set. The construction we present in proving Theorem 1.1 follows their approach, using the graph-theoretic result of [CJMS23] to obtain the same logarithmic improvement.
To prove Theorem 1.1, we will first randomly construct a discrete set $X$ of points from which to choose the centers of our copies of $K$. We will then use the graph-theoretic result of [CJMS23] to show that a large subset of $X$ may be chosen to be centers of a packing. To satisfy the preconditions of this graph-theoretic result, we will need to ensure that the copies of $K$ centered at points in $X$ do not have very large pairwise intersection. This requires some understanding of the geometry of $K$. We achieve this understanding by relating our geometric problem to Bourgain’s slicing problem. It is the precise connection between packing problems and the slicing problem that constitutes the main novelty of this work.
1.1. The Bourgain slicing problem
We now introduce the Bourgain slicing problem and survey the recent progress towards its resolution.
Definition 1.3.
Given a convex body $K$ of volume $1$, define
$L_{K}=\left(\sup_{H\subset\mathbb{R}^{d}}\operatorname{vol}(K\cap H)\right)^{-1},$
where the supremum runs over all hyperplanes $H$ passing through the origin.
Conjecture 1.4 ([Bou86, Bou87]).
There exists an absolute constant $C$ so that $L_{K}<C$, irrespective of the dimension of $K$.
While Bourgain’s conjecture has not yet been proven, there has been a flurry of recent progress. For $K\subset\mathbb{R}^{d}$, it is not too hard to prove $L_{K}\lesssim\sqrt{d}$. (See, for example, [KM22, pg. 210].) A bound of $L_{K}\lesssim d^{1/4}\log d$ was supplied by Bourgain [Bou91], shortly after the introduction of the problem. This was improved by Klartag [Kla06] to $L_{K}\lesssim d^{1/4}$.
Klartag’s bound stood for fifteen years until a breakthrough result of Chen [Che21], who gave an upper bound of $L_{K}\lesssim d^{o(1)}$. Following Chen’s result, Klartag and Lehec [KL22] proved a polylogarithmic upper bound for $L_{K}$, which was subsequently improved by Jambulapati, Lee, and Vempala [JLV22]. The strongest bound known currently is due to Klartag:
Theorem 1.5 ([Kla23]).
There exists an absolute constant $C_{\mathrm{slice}}>0$ such that, for any dimension $d\geq 1$ and any convex body $K\subset\mathbb{R}^{d}$ of volume $1$,
$L_{K}\leq C_{\mathrm{slice}}\sqrt{\log d}.$
We will use Theorem 1.5 in proving Theorem 1.1; the relationship between the two is detailed in Section 4.
Remark 1.6.
Any bound of the form $L_{K}\lesssim d^{o(1)}$ (as first shown by Chen [Che21]) is strong enough to give us our main result, albeit with a more slowly decaying $o(1)$ term than we get if we use Theorem 1.5. Any constant or polylogarithmic upper bound on $L_{K}$ (like those given by [KL22, JLV22, Kla23], or that which would hold under 1.4) yields a bound of the form $\Theta(\frac{\log\log d}{\log d})$ for the $o(1)$ term in Theorem 1.1. This is the same error that the proof in [CJMS23] achieves for the $\ell_{2}$ unit ball. If we were to use the weaker bound of $L_{K}\lesssim d^{1/4}$ or $L_{K}\lesssim d^{1/4}\log d$ (as given by [Bou91, Kla06]), we would attain
$\delta_{T}(K)\geq(1-o(1))\frac{d\log d}{2^{d+2}},$
off by a factor of two from Theorem 1.1. Correspondingly, a bound of $L_{K}\lesssim\sqrt{d}$ would not be enough to give us any nontrivial lower bound on $\delta_{T}(K)$.
It is also worth mentioning the rich connections between Bourgain’s slicing conjecture and other problems in the area of convex geometry. One particular example is the Kannan–Lovász–Simonovits conjecture [KLS95], which asks for a constant lower bound on the isoperimetric (Cheeger) constant of a $d$-dimensional isotropic log-concave distribution. (It is in terms of the KLS conjecture that Chen’s bound is phrased; see [KM22] for a discussion of the relation between the two problems.) Another is the thin-shell problem of Eldan and Klartag [EK11], which asks about the concentration of the magnitude of a vector chosen uniformly at random from such a distribution. More examples of equivalent problems are listed in [MP89, Section 5], and more detail on the historical context of and progress towards 1.4 is given in the survey article [KM22].
1.2. Special cases
Some particular choices of $K$ beside the $\ell_{2}$ unit ball warrant special mention.
$\ell_{p}$-balls
Outside of the Euclidean norm, the most natural and well-studied norms on $\mathbb{R}^{d}$ are the $\ell_{p}$ norms for $1\leq p\leq\infty$. For each fixed $p>2$, packings of density exponentially greater than $2^{-d}$ (i.e. of density at least $(2-\varepsilon_{p})^{-d}$ for some $\varepsilon_{p}>0$) were constructed in 1991 by Elkies, Odlyzko and Rush [EOR91], and improved constructions were presented by Liu and Xing [LX08]. In contrast, no such bound is known for any $p\in[1,2]$. Recently, Xie and Ge [XG22] gave two new proofs of lower bounds of the form $cd2^{-d}$ for sphere packing in $\ell_{p}$ spaces with $1<p\leq 2$ (and some generalizations thereof) with techniques from statistical physics, based on the argument of Jenssen, Joos, and Perkins [JJP19] in the $\ell_{2}$ case. Theorem 1.1 improves this bound by a logarithmic factor.
Regular simplex
The regular $(d+1)$-simplex $\triangle_{d}\subset\mathbb{R}^{d}$ is an attractive example of a convex shape about which to ask packing problems, as one of the only regular polytopes in large dimensions (along with the cross-polytope and the cube, which are $\ell_{p}$-balls for $p=1$ and $p=\infty$, respectively). Simplices do not pack translationally very tightly. Rogers and Shephard [RS57] give bounds of the form
$\frac{c_{1}d^{1/2}}{2^{2d}}\leq\delta_{T}(\triangle_{d})\leq\frac{c_{2}d^{1/2}}{2^{d}}$
for constants $c_{1},c_{2}$. Using Schmidt’s bound for packing centrally symmetric shapes, the exponent $d^{1/2}$ in the lower bound can be easily improved to $d^{3/2}$; Corollary 1.2 improves this bound by a logarithmic factor. To our knowledge, the upper bound has not been improved.
2. Outline and proof sketch
In this section, we outline our construction and state the main ingredients necessary for the proof of Theorem 1.1. We give a proof of Theorem 1.1 assuming these intermediate results. Finally, we deduce Corollary 1.2 from Theorem 1.1.
2.1. Construction of the packing
As in [CJMS23], the packings we construct will be “amorphous,” with fairly little structure. Following the argument in [CJMS23], we construct the packing in the following way, depending on a parameter $\Delta$:
  1. (1)
    Choose a set $\mathbf{X}$ of points via a Poisson point process with intensity $2^{-d}\Delta$. Construct a graph $G$ whose vertex set is $\mathbf{X}$ and whose edges are pairs $(x,y)$ for which $x+K$ and $y+K$ intersect. (Vertices in this graph should have degree approximately $\Delta$.)
  2. (2)
    Remove from $\mathbf{X}$ those vertices which have degree much larger than $\Delta$, or which have codegree more than $d^{-9}\Delta$ with some other vertex of $\mathbf{X}$. Call the resulting graph $G^{\prime}$.
  3. (3)
    Find a large independent set $A\subset X$ in $G^{\prime}$. The points corresponding to the vertices in $A$ form the centers of the copies of $K$ in our packing.
The resulting packing will have density approximately $2^{-d}\log\Delta$, improving on the previous bound $cd2^{-d}$ as long as $\Delta$ can be taken to be super-exponential in $d$.
2.2. Structure of the proof
Let $K\subset\mathbb{R}^{d}$ be a compact, centrally symmetric convex set of volume $1$. Define the set
$I_{K}=\left\{x\in\mathbb{R}^{d}\colon\operatorname{vol}(K\cap(K+x))>d^{-10}\right\},$
and select $\Delta_{K}=(d\operatorname{vol}(I_{K}))^{-1}$. Given a finite set of points $X$, let $G(X,K)$ be the graph with vertex set $X$ and with an edge $xy$ if and only if $x-y\in 2K$, i.e. if and only if the translates $x+K$ and $y+K$ intersect. The first and second steps in the above outline are accomplished via the following lemma.
Lemma 2.1 (cf. [CJMS23, Lemma 2.1]).
Suppose $d>10$ and $\Delta$ is so that $d^{12}<\Delta\leq\Delta_{K}$. Let $\Omega\subset\mathbb{R}^{d}$ be bounded and measurable. There exists some finite set of points $X\subset\Omega$ such that
  • $|X|\geq(1-2/d)(\Delta/2^{d})\operatorname{vol}(\Omega)$,
  • the graph $G(X,K)$ has maximum degree at most $\Delta+\Delta^{2/3}$, and
  • the graph $G(X,K)$ has maximum codegree at most $d^{-9}\Delta$.
(Here, the codegree of a pair of distinct vertices $(v_{1},v_{2})$ is the number of vertices $w$ of the graph for which both $v_{1}w$ and $v_{2}w$ are edges.) We prove Lemma 2.1 in Section 3, following the above outline. To accomplish the third step of the construction, we use the following graph-theoretic result.
Theorem 2.2 ([CJMS23, Theorem 1.3]).
Let $G$ be a graph on $n$ vertices with maximum degree at most $\Delta$, and suppose that the maximum codegree of any two vertices in $G$ is at most $\Delta/(2\log\Delta)^{7}$. Then there exists an independent set in $G$ of size at least
$(1-o(1))\frac{n\log\Delta}{\Delta},$
where the $o(1)$ tends to zero as $\Delta\to\infty$.
To complete the proof of Theorem 1.1, we must lower-bound $\Delta_{K}$; equivalently, we must give an upper bound on the volume of $I_{K}$. Indeed, the following holds.
Proposition 2.3.
There exists a constant $C_{\mathrm{int}}>0$ for which, for any centrally symmetric convex set $K\subset\mathbb{R}^{d}$ of volume $1$,
$\operatorname{vol}(I_{K})\leq\left(\frac{C_{\mathrm{int}}\log^{3}d}{d}\right)^{d/2}.$
It is in proving Proposition 2.3 that we will use bounds on $L_{K}$, i.e. Theorem 1.5. See Section 4 for a proof of Proposition 2.3 and more discussion of the connection between $I_{K}$ and Bourgain’s slicing problem.
Proof of Theorem 1.1.
For each compact $\Omega\subset\mathbb{R}^{d}$, we find a finite set $A\subset\Omega$ so that
$|A|\geq(1-o(1))\frac{d\log d}{2^{d+1}}\operatorname{vol}(\Omega)$
and the balls $x+K$ and $y+K$ are disjoint for $x,y\in A$. Via a standard compactness argument (see [BMP05, Section 1.1]), having such a packing within each compact set is enough to give the desired bound on $\delta_{T}(K)$.
We may assume $d$ is sufficiently large: for $d$ small, we can use the trivial bound $\delta_{T}(K)\geq 2^{-d}$ and absorb the loss into the $o(1)$ term. Let $\Delta=\min\big{(}\Delta_{K},d^{d^{8/7}}\big{)}$, so that (using Proposition 2.3) we have $\Delta_{K}\geq\Delta>d^{12}$. So, by Lemma 2.1, we may choose a set $X\subset\Omega$ with
$|X|\geq\frac{(1-o(1))\Delta\operatorname{vol}(\Omega)}{2^{d}}$
so that $G(X,K)$ has maximum degree at most $\Delta+\Delta^{2/3}$ and maximum codegree at most $d^{-9}\Delta$. Since $\Delta\leq d^{d^{8/7}}$, we have $d^{-9}\Delta\leq\Delta/(2\log\Delta)^{7}$. So, we may apply Theorem 2.2 to the graph $G(X,K)$ to find an independent set $A\subset X$ with
$|A|\geq(1-o(1))\frac{|X|\log\Delta}{\Delta}\geq\frac{(1-o(1))\operatorname{vol}(\Omega)\log\Delta}{2^{d}}.$
This set $A$ is our set of centers. Finally, Proposition 2.3 furnishes that
$\log\Delta$ $\geq\min\left(d^{8/7}\log d,\log\Delta_{K}\right)$
$=\min\left(d^{8/7}\log d,\log\left((d\operatorname{vol}I_{K})^{-1}\right)\right)$
$\geq\frac{(1-o(1))d\log d}{2}.$
This finishes the proof of Theorem 1.1. ∎
2.3. Non-symmetric sets
Here, we explain how to deduce Corollary 1.2 from Theorem 1.1. This procedure goes back to Rogers and Shephard [RS57], who coupled their bound on the volume of $K-K$ with an observation of Minkowski [Min04] to attain the bound
$\delta_{T}(K)\geq(1-o(1))\frac{\sqrt{\pi d}}{2^{2d}}$
for any compact convex set $K$. For completeness, we outline the deduction.
Lemma 2.4 ([Min04]).
For any convex body $K\subset\mathbb{R}^{d}$,
$\frac{\delta_{T}(K)}{\operatorname{vol}K}=\frac{\delta_{T}\left(\frac{K-K}{2}\right)}{\operatorname{vol}\left(\frac{K-K}{2}\right)}.$
Proof sketch.
This inequality follows from the fact that a discrete set $A\subset\mathbb{R}^{d}$ serves as the centers for a packing of $K$ if and only if it serves as the centers for a packing of $\frac{K-K}{2}$. ∎
Lemma 2.5 (Rogers–Shephard inequality; [RS57]).
For any convex body $K\subset\mathbb{R}^{d}$,
$\operatorname{vol}(K-K)\leq\binom{2d}{d}\operatorname{vol}K,$
with equality if and only if $K$ is a simplex.
We remark that, since equality in Lemma 2.5 holds in the case of a simplex $K=\triangle_{d}$, the trivial bound $\delta_{T}(\frac{\triangle_{d}-\triangle_{d}}{2})\leq 1$ furnishes the upper bound $\delta_{T}(\triangle_{d})\leq 2^{d}/\binom{2d}{d}\sim 2^{-d}\sqrt{\pi d}$ mentioned in the introduction.
Proof of Corollary 1.2.
Note that $L:=\frac{K-K}{2}$ is convex and centrally symmetric. So, chaining Lemma 2.4, Lemma 2.5, and Theorem 1.1 gives
$\delta_{T}(K)=\frac{\operatorname{vol}K}{\operatorname{vol}L}\delta_{T}(L)\leq\frac{2^{d}}{\binom{2d}{d}}\delta_{T}(L)\leq\frac{(1-o(1))d\log d}{2\binom{2d}{d}}.$
The corollary then follows from the asymptotic $\binom{2d}{d}=2^{2d}(\pi d)^{-1/2}(1+o(1))$. ∎
3. Selecting the set $X$: Proof of Lemma 2.1
We now prove Lemma 2.1. As previously mentioned, our strategy will be to first pick a set of points $\mathbf{X}\subset\Omega$ via a Poisson point process and then to remove a $o(1)$-fraction of the points to form $X$. The proof is quite similar to the proof of [CJMS23, Lemma 2.1]. The additional complexity is in bounding the maximum codegree, and this is where the choice of $\Delta_{K}$ as determined by $\operatorname{vol}(I_{K})$ is relevant. We will need the following standard tail bound for Poisson random variables.
Lemma 3.1 (Follows from [JŁR00, Theorem 2.1]).
A random variable $Z\sim\operatorname{Pois}(\lambda)$ satisfies the following upper tail bound: for each $t\geq 1$,
$\Pr\big{[}Z>(1+t)\lambda\big{]}\leq e^{-\lambda t/3}.$
We now show Lemma 2.1.
Proof of Lemma 2.1.
Through what follows, we write $I=I_{K}$, as the relevant set $K$ will not change. We choose the set $X$ in the following manner. Let $\mathbf{X}$ be a set of points chosen according to a Poisson point process inside $\Omega$ with intensity $\lambda:=2^{-d}\Delta$. Define the following subsets $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\subset\mathbf{X}$:
$\mathbf{X}_{1}$ =$=\left\{x\in\mathbf{X}\colon|\mathbf{X}\cap(x+2K)|>\Delta+\Delta^{2/3}\right\},$
$\mathbf{X}_{2}$ =$=\left\{x\in\mathbf{X}\colon\mathbf{X}\cap(x+2I)\text{ contains apoint other than $x$}\right\},$
$S_{3}$ =$=\left\{(x,y)\in\mathbf{X}^{2}\colon x-y\not\in 2I\text{ and }|(\mathbf{X}\setminus\{x,y\})\cap(x+2K)\cap(y+2K)|\geq d^{-9}\Delta\right\},$
$\mathbf{X}_{3}$ =$=\left\{x\in\mathbf{X}\colon\exists y\in\mathbf{X}\text{ with }(x,y)\in S_{3}\right\}.$
Set $X=\mathbf{X}\setminus(\mathbf{X}_{1}\cup\mathbf{X}_{2}\cup\mathbf{X}_{3})$. We make the following observations.
  • By the definition of $\mathbf{X}_{1}$, the degree of any vertex of $G(X,K)$ is at most $\Delta+\Delta^{2/3}$.
  • By the definition of $\mathbf{X}_{2}$, each pair $(x,y)$ of distinct elements of $X$ satisfies $x-y\not\in 2I$.
  • By the definition of $\mathbf{X}_{3}$ and the above property, the maximum codegree of $G(X,K)$ is at most $d^{-9}\Delta$.
So, as long as $|X|\geq(1-2/d)(\Delta/2^{d})\operatorname{vol}(\Omega)$, $X$ will satisfy the conditions of the lemma. We will in fact show that $\mathbb{E}|X|$ exceeds this quantity; the result follows by Markov’s inequality. We have
$\mathbb{E}|\mathbf{X}|$ $=\lambda\operatorname{vol}(\Omega)=\Delta/2^{d}\cdot\operatorname{vol}(\Omega)\text{ and }$
$\mathbb{E}|X|$ $\geq\mathbb{E}|\mathbf{X}|-\mathbb{E}|\mathbf{X}_{1}|-\mathbb{E}|\mathbf{X}_{2}|-\mathbb{E}|S_{3}|.$
Hence, it suffices to show the following bounds:
(3.1) $\mathbb{E}|\mathbf{X}_{1}|$ $\leq e^{-d}\mathbb{E}|\mathbf{X}|$
(3.2) $\mathbb{E}|\mathbf{X}_{2}|$ $\leq d^{-1}\mathbb{E}|\mathbf{X}|$
(3.3) $\mathbb{E}|S_{3}|$ $\leq e^{-d}\mathbb{E}|\mathbf{X}|.$
Proof of (3.1).
Conditioned on $x\in\mathbf{X}$, $|\mathbf{X}\cap(x+2K))|-1$ is Poisson with mean
$\lambda\operatorname{vol}\big{(}(x+2K)\cap\Omega\big{)}\leq 2^{d}\lambda=\Delta.$
So, Lemma 3.1 implies
$\Pr[x\in\mathbf{X}_{1}|x\in\mathbf{X}]\leq e^{-(\Delta^{2/3}-1)/3}.$
This bound, together with the fact that $\Delta>d^{12}$, gives (3.1).
Proof of (3.2).
Conditioned on $x\in\mathbf{X}$, the random variable $|\mathbf{X}\cap(x+2I)|-1$ is Poisson with mean
$\lambda\operatorname{vol}\big{(}(x+2I)\cap\Omega\big{)}\leq\lambda\operatorname{vol}(2I)=\Delta\operatorname{vol}(2I)2^{-d}=\Delta\operatorname{vol}(I)\leq\frac{1}{d}$
So, $|\mathbf{X}\cap(x+2I)|>1$, i.e. $x\in\mathbf{X}_{2}$, with probability at most $1-e^{-1/d}\leq 1/d$. This gives (3.2).
Proof of (3.3).
If $(x,y)\in S_{3}$, then $x+2K$ and $y+2K$ must intersect, so $x-y\in 4K$. We now show that, for any pair of distinct points $(x,y)$ with $x-y\in 4K$,
$\Pr\left[(x,y)\in S_{3}\big{|}x,y\in\mathbf{X}\right]\leq e^{-d^{-10}\Delta}.$
Indeed, fixing $x$ and $y$ with $x-y\not\in 2I$ and conditioning on $x,y\in\mathbf{X}$, we have
$\big{|}(\mathbf{X}\setminus\{x,y\})\cap(x+2K)\cap(y+2K)\big{|}\sim\operatorname{Pois}\big{(}\lambda\operatorname{vol}((x+2K)\cap(y+2K))\big{)},$
and
$\operatorname{vol}\big{(}(x+2K)\cap(y+2K)\big{)}=2^{d}\operatorname{vol}\left(K\cap\left(K+\frac{x-y}{2}\right)\right)\leq 2^{d}d^{-10}$
since $(x-y)/2\not\in I$. We conclude that $|(\mathbf{X}\setminus\{x,y\})\cap(x+2K)\cap(y+2K)|$ is Poisson with mean at most $d^{-10}\Delta$. As a result, the probability that it exceeds $d^{-9}\Delta$ is at most, by Lemma 3.1,
$\Pr\left[\operatorname{Pois}(d^{-10}\Delta)\geq 4d^{-10}\Delta\right]\leq e^{-d^{-10}\Delta},$
as desired. Now, we can use this bound to upper-bound $\mathbb{E}|S_{3}|$:
$\mathbb{E}|S_{3}|$ $\leq e^{-d^{-10}\Delta}\cdot\mathbb{E}\big{|}\{(x,y)\colon x,y\in\mathbf{X},\ x\neq y,\ x-y\in 4K\}\big{|}$
$\leq e^{-d^{-10}\Delta}\cdot\mathbb{E}|\mathbf{X}|\cdot\sup_{x\in\Omega}\mathbb{E}\left[\big{|}\{y\in\mathbf{X}\setminus\{x\}\colon x-y\in 4K\}\big{|}|x\in\mathbf{X}\right]$
$=e^{-d^{-10}\Delta}\cdot\mathbb{E}|\mathbf{X}|\cdot\lambda\operatorname{vol}(4K)=2^{d}\Delta e^{-d^{-10}\Delta}\cdot\mathbb{E}|\mathbf{X}|.$
The fact that $\Delta>d^{12}$ is enough to give (3.3).
4. Bounding the volume of $I_{K}$: Proof of Proposition 2.3
We now prove Proposition 2.3, i.e. that there exists a constant $C_{\mathrm{int}}>0$ so that, for every dimension $d$ and every centrally symmetric convex body $K$
$\operatorname{vol}(I_{K})\leq\left(\frac{C_{\mathrm{int}}\log^{3}d}{d}\right)^{d/2},$
where as before
$I_{K}=\left\{x\in\mathbb{R}^{d}\colon\operatorname{vol}(K\cap(K+x))>d^{-10}\right\}.$
As previously mentioned, we will need to use results pertaining to the Bourgain slicing problem. We begin by describing how an upper bound on $L_{K}$ can be used to control intersections of arbitrary convex sets with arbitrary hyperplanes.
4.1. Intersecting convex bodies with hyperplanes
Our first step in proving Proposition 2.3 will be to apply a linear transformation to transform $K$ into a more agreeable form.
Definition 4.1.
A convex body $K\subset\mathbb{R}^{d}$ of volume $1$ is in isotropic position if there exists some real $\alpha$ so that
$\mathbb{E}_{Y\sim\operatorname{Unif}(K)}Y=0\text{\quad and\quad}\mathbb{E}_{Y\sim\operatorname{Unif}(K)}YY^{\intercal}=\alpha I_{d},$
where $I_{d}$ is the $d\times d$ identity matrix.
Every convex body $K\subset\mathbb{R}^{d}$ of volume $1$ can be put in isotropic position by a volume-preserving affine transformation. Indeed, first translate $K$ to have barycenter 00. Then, the matrix $\mathbb{E}_{Y\sim\operatorname{Unif}(K)}YY^{\intercal}$ is positive definite, and so it can be written as $AA^{\intercal}$ for some invertible matrix $A$. When $K$ is scaled by an appropriate multiple of $A^{-1}$, the result is in isotropic position.
An important aspect of isotropic position, with a view to apply Theorem 1.5, will be that the volumes of different hyperplane sections of $K$ cannot differ too much.
Lemma 4.2 ([Hen80, Fra99]).
There exists some absolute constant $C_{\mathrm{ratio}}$ so that the following holds: If $K$ is a convex body in isotropic position and $H_{1}$ and $H_{2}$ are hyperplanes passing through the origin, then
$\frac{\operatorname{vol}(K\cap H_{1})}{\operatorname{vol}(K\cap H_{2})}\leq C_{\mathrm{ratio}}.$
In the case where $K$ is centrally symmetric, which is all that is needed for our purposes, this result was proven by Hensley [Hen80]. Lemma 4.2 was proven for general $K$ by Fradelizi [Fra99]; in fact, Fradelizi showed that the optimal constant is $C_{\mathrm{ratio}}=\sqrt{6}$.
4.2. Log-concave probability distributions
Some key tools in our proof come from the theory of log-concave probability distributions.
Definition 4.3.
A probability distribution $\mu$ on $\mathbb{R}^{d}$ is log-concave if its support is convex and, for any compact $A,B\subset\mathbb{R}^{d}$ and any $10<\lambda<$ 1,
$\mu(\lambda A+(1-\lambda)B)\geq\mu(A)^{\lambda}\mu(B)^{1-\lambda}.$
The Brunn–Minkowski inequality implies that the uniform measure on any convex set $K$ is log-concave. We will need the following two facts about log-concave distributions. The first follows from the Prékopa–Leindler inequality, and is important for reducing questions about high-dimensional convex sets down to one dimension. See [Pré73] for more discussion.
Lemma 4.4 ([Pré73, Theorem 8]).
For any log-concave distribution $\mu$ on $\mathbb{R}^{d}$ and any projection $\pi:\mathbb{R}^{d}\to\mathbb{R}^{k}$, the pushforward $\pi_{*}\mu$ is also log-concave.
The second fact is an exponentially decaying tail bound for one-dimensional log-concave distributions $\mu$. It will be useful for us to control precisely how the tails depend on the density function of $\mu$, so we state and prove a version tailored to our situation.
Lemma 4.5.
Let $\mu$ be any symmetric log-concave probability measure on $\mathbb{R}$ with density function $p_{\mu}$. We have, for any $t>0$,
$\Pr_{x\sim\mu}[x>t]\leq\frac{1}{\log 2}2^{-tp_{\mu}(0)}.$
Proof.
Let $v=p_{\mu}(0)^{-1}$. First, we note that $p_{\mu}(v)\leq 1/(2v)$, as otherwise $p_{\mu}(x)>1/(2v)$ for each $x\in[-v,v]$, which contradicts the fact that $p_{\mu}$ is a probability measure. Now, for $x>v$, the log-concavity of $\mu$ gives
$p_{\mu}(x)\leq p_{\mu}(0)\left(\frac{p_{\mu}(v)}{p_{\mu}(0)}\right)^{x/v}\leq\frac{1}{v2^{x/v}}.$
For $t<v$, the result follows from the fact that $\Pr_{x\sim\mu}[x>0]=1/2$. Otherwise, the above gives
[…]
as desired. ∎
4.3. Proof of Proposition 2.3
We are now ready to prove Proposition 2.3. To explicate the dependence of $\operatorname{vol}(I_{K})$ on $L_{K}$, we begin by proving the following:
Lemma 4.6.
There exists some absolute constant $C_{\mathrm{vol}}>0$ so that, for any symmetric isotropic convex body $K\subset\mathbb{R}^{d}$ of volume $1$, we have
$\operatorname{vol}I_{K}\leq\left(\frac{C_{\mathrm{vol}}L_{K}\log d}{\sqrt{d}}\right)^{d}.$
Proof.
In fact, we can take $C_{\mathrm{vol}}=240\cdot C_{\mathrm{ratio}}$.
We let $\lVert\cdot\rVert=\lVert\cdot\rVert_{2}$ be the standard Euclidean norm on $\mathbb{R}^{d}$. The main claim is that, for any $x\in\mathbb{R}^{d}$,
$\operatorname{vol}\big{(}(K-x)\cap(K+x)\big{)}\leq 3\exp\left(-\frac{\|x\|}{2C_{\mathrm{ratio}}L_{K}}\right).$
If $x=0$, this inequality is clear. Otherwise, let $\hat{x}=x/\|x\|$. Let $\pi:\mathbb{R}^{d}\to\mathbb{R}$ be the projection $y\mapsto\hat{x}\cdot y$, let $Y\sim\operatorname{Unif}(K)$ be a random variable, and let $p_{x}:\mathbb{R}\to[0,\infty)$ be the probability density function of $\pi(Y)$. Since $\operatorname{vol}K=1$, $p_{x}$ gives the volume of each “slice” of $K$:
$p_{x}(t)=\operatorname{vol}(K\cap\pi^{-1}(t)).$
We now bound the volume of the slices of $(K-x)\cap(K+x)$. For each real $t\geq 0$, we compute
$(K-x)\cap(K+x)\cap\pi^{-1}(t)$ $\subset(K-x)\cap\pi^{-1}(t)$
$=-x+\big{(}K\cap(\pi^{-1}(t)+x)\big{)}$
$=-x+\big{(}K\cap\pi^{-1}(t+\|x\|)\big{)},$
and so
$\operatorname{vol}\big{(}(K-x)\cap(K+x)\cap\pi^{-1}(t)\big{)}\leq\operatorname{vol}\big{(}K\cap\pi^{-1}(t+\|x\|)\big{)}=p_{x}(t+\|x\|).$
For $t\leq 0$, we compute the volume of the intersection $(K+x)\cap\pi^{-1}(t)$ instead, and get
$\operatorname{vol}\big{(}(K-x)\cap(K+x)\cap\pi^{-1}(t)\big{)}\leq\operatorname{vol}\big{(}K\cap\pi^{-1}(t-\|x\|)\big{)}=p_{x}(t-\|x\|).$
As a result,
[…]
$=\Pr\big{[}\pi(Y)\leq-\|x\|\big{]}+\Pr\big{[}\pi(Y)\geq\|x\|\big{]}.$
Now, the random variable $Y$ is log-concave. So, by Lemma 4.4, $\pi(Y)$ is also log-concave. We may lower-bound $p_{x}(0)$ using Lemma 4.2:
$p_{x}(0)=\operatorname{vol}(K\cap\pi^{-1}(0))$ $\geq\inf_{H}\operatorname{vol}(K\cap H)$
$\geq\frac{1}{C_{\mathrm{ratio}}}\sup_{H}\operatorname{vol}(K\cap H)=\frac{1}{C_{\mathrm{ratio}}L_{K}},$
where both $\inf_{H}$ and $\sup_{H}$ are over hyperplanes passing through the origin. Lemma 4.5 now gives
$\operatorname{vol}\big{(}(K-x)\cap(K+x)\big{)}$ $\leq\Pr_{Y\sim\operatorname{Unif}(K)}\big{[}|\pi(Y)|\geq\|x\|\big{]}$
$\leq\frac{2}{\log 2}\cdot 2^{-\|x\|p_{x}(0)}$
$\leq 3\exp\left(-\frac{\|x\|}{2C_{\mathrm{ratio}}L_{K}}\right)$
as claimed. We conclude that, for any $x\in I_{K}$,
$d^{-10}\leq\operatorname{vol}\big{(}K\cap(K+x)\big{)}=\operatorname{vol}\left(\left(K-\frac{1}{2}x\right)\cap\left(K+\frac{1}{2}x\right)\right)\leq 3\exp\left(-\frac{\|x\|}{4C_{\mathrm{ratio}}L_{K}}\right).$
Rearranging the above, we get that $\|x\|\leq 48C_{\mathrm{ratio}}L_{K}\log d$ for all $x\in I_{K}$. Thus $I_{K}$ is contained within the (Euclidean) ball of radius $48C_{\mathrm{ratio}}L_{K}\log d$ centered at zero. This containment gives
$\operatorname{vol}I_{K}\leq\frac{\pi^{d/2}(48C_{\mathrm{ratio}}L_{K}\log d)^{d}}{(d/2)!}\leq\left(\frac{240C_{\mathrm{ratio}}L_{K}\log d}{\sqrt{d}}\right)^{d}.$
Setting $C_{\mathrm{vol}}=240C_{\mathrm{ratio}}$ concludes the proof. ∎
Combining Lemma 4.6 and Theorem 1.5, we get Proposition 2.3.
Proof of Proposition 2.3.
We can take $C_{\mathrm{int}}=(C_{\mathrm{vol}}\cdot C_{\mathrm{slice}})^{1/2}$. First, we reduce to the isotropic case. Let $K^{\prime}$ be the image of $K$ by a volume-preserving linear transformation $A\in\operatorname{SL}(d)$ so that $K^{\prime}$ is isotropic. Then $I_{K^{\prime}}$ is the image of $I_{K}$ by $A$. Since $K^{\prime}$ is isotropic, we may apply Lemma 4.6 to get
$\operatorname{vol}I_{K}=\operatorname{vol}I_{K^{\prime}}\leq\left(\frac{C_{\mathrm{vol}}L_{K^{\prime}}\log d}{\sqrt{d}}\right)^{d}.$
By Theorem 1.5, we have $L_{K^{\prime}}\leq C_{\mathrm{slice}}\sqrt{\log d}$. Combining this with the above gives the claimed result. ∎
Acknowledgement
The author would like to thank Jacob Fox for many helpful conversations and suggestions.