### v2

# Lower bounds for sphere packing in arbitrary norms

Abstract.
We show that in any $d$-dimensional real normed space, unit balls can be packed with density at least
$\frac{(1-o(1))d\log d}{2^{d+1}},$
improving a result of Schmidt from 1958 by a logarithmic factor and generalizing the recent result of Campos, Jenssen, Michelen, and Sahasrabudhe in the $\ell_{2}$ norm. Our main tools are the graph-theoretic result used in the $\ell_{2}$ construction and volume bounds from convex geometry due to Petty and Schmuckenschläger.
1. Introduction
The question of how densely congruent spheres may be packed in $d$-dimensional space is a classical one. The “trivial bound,” i.e. that at least a $2^{-d}$ proportion of space may be covered, was first improved by a factor of $2$ by Minkowski [Min05] in 1905. The first asymptotic improvement (i.e., by more than a constant factor) on the trivial bound was due to Rogers [Rog47], who gave a lower bound of the form $c(1-o(1))d2^{-d}$ for $c=2/e$. This constant was improved to $c=2$ in 1992 by Ball [Bal92] and to $c=65963$ by Venkatesh in 2012 [Ven13]. Very recently, Campos, Jenssen, Michelen, and Sahasrabudhe [CJMS23] gave an asymptotic improvement, replacing the constant $c$ by $\frac{1}{2}\log d$ [^1] Here and throughout, $\log$ denotes the natural logarithm.
[…]
Our main result extends the result of [CJMS23] to arbitrary norms, with the same asymptotic improvement and the same constant factor. This improves upon the previous best-known bound by a logarithmic factor. We will phrase our results in terms of centrally symmetric convex bodies (i.e., centrally symmetric compact convex sets with nonempty interior), as our arguments are more natural in this language. [^2] Convex bodies symmetric about 00 and norms on $\mathbb{R}^{d}$ are in bijection: the unit ball $\{x\in\mathbb{R}^{d}:\|x\|\leq 1\}$ has nonempty interior, is centrally symmetric, is compact, and is convex, while if $K$ has nonempty interior and is centrally symmetric, compact, and convex, then $\|x\|=\inf\{\lambda\geq 0:x\in\lambda K\}$ is a norm on $\mathbb{R}^{d}$. Given a compact set $K\subset\mathbb{R}^{d}$ with nonempty interior, let $\delta_{T}(K)$ be the maximum density of a packing of translates of $K$ in $\mathbb{R}^{d}$.
Theorem 1.1.
For each compact, centrally symmetric convex body $K\subset\mathbb{R}^{d}$,
$\delta_{T}(K)\geq(1-o(1))\frac{d\log d}{2^{d+1}},$
where the $o(1)$ term tends to zero with the dimension $d$, irrespective of the particular body $K$.
We quickly mention the question of finding upper bounds. In the $\ell_{2}$ case, the best upper bound is of the shape $2^{-(0.599\ldots-o(1))d}$, due to Kabatjanskiĭ and Levenšteĭn [KL78] with constant-factor improvements by Cohn and Zhao [CZ14] and Sardari and Zargar [SZ23]. Some bodies $K$ (such as the hypercube) can of course be packed much more tightly, so there is no general-purpose upper bound in the setting of Theorem 1.1. We are not aware of any centrally symmetric convex body $K\subset\mathbb{R}^{d}$ whose translational packing density is known to be strictly less than the best known upper bound for the $\ell_{2}$ unit ball.
It is also interesting to consider translational packings of shapes which are not centrally symmetric. Via a result of Rogers and Shephard [RS57], we may the following from Theorem 1.1.
Corollary 1.2.
For each compact, convex body $K\subset\mathbb{R}^{d}$,
$\delta_{T}(K)\geq(1-o(1))\frac{\sqrt{\pi}d^{3/2}\log d}{2^{2d+1}}.$
We now make some comments about our approach. Historically, most of the arguments achieving dense packings of convex shapes have constructed lattice packings, i.e. packings in which the set of centers forms a lattice in $\mathbb{R}^{d}$. In contrast, the packing given in [CJMS23] is amorphous, lacking any global structure. In the case of the $\ell_{2}$ unit ball, densities asymptotic to $cd2^{-d}$ were first achieved for amorphous packings in 2004 by Krivelevich, Litsyn, and Vardy [KLV04], with constant-factor improvements by Jenssen, Joos, and Perkins [JJP19] and Fernández, Kim, Liu, and Pikhurko [FKLP23]. For $\ell_{p}$ unit balls and some generalizations (superballs), Xie and Ge [XG22] recently extended the techniques of [JJP19] to give lower bounds of the form $cd2^{-d}$. To our knowledge, ours is the first work achieving densities asymptotically greater than $2^{-d}$ for other convex bodies using amorphous packings.
In [KLV04], the analysis of the amorphous packing is via tools from graph theory. Specifically, the authors use a result of Ajtai, Komlós, and Szemerédi [AKS80] that a graph with few triangles must have a large independent set. The improvement of Campos, Jenssen, Michelen, and Sahasrabudhe comes from a novel and more refined graph-theoretic result, which allows them to obtain an extra factor of $\log d$ in the size of the independent set. The construction we present in proving Theorem 1.1 follows their approach, using the graph-theoretic result of [CJMS23] to obtain the same logarithmic improvement.
To prove Theorem 1.1, we will first randomly construct a discrete set $X$ of points from which to choose the centers of our copies of $K$. We will then use the graph-theoretic result of [CJMS23] to show that a large subset of $X$ may be chosen to be centers of a packing. To satisfy the preconditions of this graph-theoretic result, we will need to ensure that the copies of $K$ centered at points in $X$ do not have very large pairwise intersection. This requires some understanding of the geometry of $K$. We achieve this understanding by considering the polar projection body of $K$, a centrally symmetric convex body associated to $K$ (see Definition 4.1). A result of Schmuckenschläger [Sch92] reduces our problem to studying the volume of the polar projection body; to do this, we appeal to Petty’s projection inequality [Pet71]. This argument is carried out in Section 4.
In an earlier version of this article, the geometric argument proceeded instead via a reduction to the Bourgain slicing problem. From this reduction, we attained the main result by appealing to recent work [Che21, KL22, JLV22, Kla23, Gua24, KL24] on the slicing problem. This approach is sketched in Remark 4.4.
1.1. Special cases
Some particular choices of $K$ beside the $\ell_{2}$ unit ball warrant special mention.
$\ell_{p}$-balls
Outside of the Euclidean norm, the most natural and well-studied norms on $\mathbb{R}^{d}$ are the $\ell_{p}$ norms for $1\leq p\leq\infty$. For each fixed $p>2$, packings of density exponentially greater than $2^{-d}$ (i.e. of density at least $(2-\varepsilon_{p})^{-d}$ for some $\varepsilon_{p}>0$) were constructed in 1991 by Elkies, Odlyzko and Rush [EOR91], and improved constructions were presented by Liu and Xing [LX08]. In contrast, no such bound is known for any $p\in[1,2]$. Recently, Xie and Ge [XG22] gave two new proofs of lower bounds of the form $cd2^{-d}$ for sphere packing in $\ell_{p}$ spaces with $1<p\leq 2$ (and some generalizations thereof) with techniques from statistical physics, based on the argument of Jenssen, Joos, and Perkins [JJP19] in the $\ell_{2}$ case. Theorem 1.1 improves this bound by a logarithmic factor.
Regular simplex
The regular $d$-simplex $\triangle_{d}\subset\mathbb{R}^{d}$ is an attractive example of a convex shape about which to ask packing problems, as one of the only regular polytopes in large dimensions (along with the cross-polytope and the cube, which are $\ell_{p}$-balls for $p=1$ and $p=\infty$, respectively). Simplices do not pack translationally very tightly. Rogers and Shephard [RS57] give bounds of the form
(1.1) $\frac{c_{1}d^{1/2}}{2^{2d}}\leq\delta_{T}(\triangle_{d})\leq\frac{c_{2}d^{1/2}}{2^{d}}$
for constants $c_{1},c_{2}$. Using Schmidt’s bound for packing centrally symmetric shapes, the exponent $d^{1/2}$ in the lower bound of (1.1) can be easily improved to $d^{3/2}$; Corollary 1.2 improves this bound by a logarithmic factor.
Remark 1.3.
The upper bound in (1.1) may be improved (for large $d$) to $2^{-c_{3}d}$ for some $c_{3}>1.428$. Let $\triangle_{d}$ be the $d$-simplex in $\mathbb{R}^{d+1}$ of side length $\sqrt{2}$ formed as the convex hull of the standard basis vectors. The key observation is that the symmetric convex body $\frac{\triangle_{d}-\triangle_{d}}{2}$ is exactly the intersection of the $(d+1)$-dimensional cross-polytope, i.e. the $\ell_{1}$ unit ball $B_{\ell^{1}}^{d+1}$ in $\mathbb{R}^{d+1}$, with the hyperplane $H$ described by $x_{1}+\cdots+x_{d+1}=0$. We sketch the remainder of the argument: if a set $A\subset H$ serves as the set of centers for a packing of $\frac{\triangle_{d}-\triangle_{d}}{2}$, then $A+\frac{2}{d+1}(1,1,\ldots,1)$ forms a set of centers for a packing of $B_{\ell^{1}}^{d+1}$. This gives the inequality
(1.2) $\delta_{T}\left(B_{\ell^{1}}^{d+1}\right)\geq\frac{\operatorname{vol}(B_{\ell^{1}}^{d+1})}{\operatorname{vol}\big{(}\frac{\triangle_{d}-\triangle_{d}}{2}\big{)}}\delta_{T}\big{(}\tfrac{\triangle_{d}-\triangle_{d}}{2}\big{)}\frac{\sqrt{d+1}}{2}\geq\frac{c}{\sqrt{d}}\delta_{T}\big{(}\tfrac{\triangle_{d}-\triangle_{d}}{2}\big{)}$
for some constant $c>0$ (the $\frac{1}{2}\sqrt{d+1}$ factor comes from the distinction between the sets of centers for the two convex bodies). A result of Rankin [Ran48] gives an upper bound $\delta_{T}(B_{\ell^{1}}^{d+1})\leq 2^{-(0.428\ldots)d}$; combining this with (1.2) gives a similar upper bound on $\delta_{T}(\frac{\triangle_{d}-\triangle_{d}}{2})$. This bound, in turn, yields an upper bound of $2^{-(1.428\ldots)d}$ on $\delta_{T}(\triangle_{d})$ using the argument of Rogers and Shephard (see Section 2.3). (We thank Dmitrii Zakharov for bringing the connection between these two convex bodies to our attention.)
2. Outline and proof sketch
In this section, we outline our construction and state the main ingredients necessary for the proof of Theorem 1.1. We give a proof of Theorem 1.1 assuming these intermediate results. Finally, we deduce Corollary 1.2 from Theorem 1.1.
2.1. Construction of the packing
As in [CJMS23], the packings we construct will be “amorphous,” with fairly little structure. Following the argument in [CJMS23], we construct the packing in the following way, depending on a parameter $\Delta$:
  1. (1)
    Choose a set $\mathbf{X}$ of points via a Poisson point process with intensity $2^{-d}\Delta$. Construct a graph $G$ whose vertex set is $\mathbf{X}$ and whose edges are pairs $(x,y)$ for which $x+K$ and $y+K$ intersect. (Vertices in this graph should have degree approximately $\Delta$.)
  2. (2)
    Remove from $\mathbf{X}$ those vertices which have degree much larger than $\Delta$, or which have codegree more than $d^{-9}\Delta$ with some other vertex of $\mathbf{X}$. Call the resulting graph $G^{\prime}$.
  3. (3)
    Find a large independent set $A\subset X$ in $G^{\prime}$. The points corresponding to the vertices in $A$ form the centers of the copies of $K$ in our packing.
The resulting packing will have density approximately $2^{-d}\log\Delta$, improving on the previous bound $cd2^{-d}$ as long as $\Delta$ can be taken to be super-exponential in $d$.
2.2. Structure of the proof
Let $K\subset\mathbb{R}^{d}$ be a centrally symmetric convex body of volume $1$. Define the set
$I_{K}=\left\{x\in\mathbb{R}^{d}:\operatorname{vol}(K\cap(K+x))>d^{-10}\right\},$
and select $\Delta_{K}=(d\operatorname{vol}(I_{K}))^{-1}$. Given a finite set of points $X$, let $G(X,K)$ be the graph with vertex set $X$ and with an edge $xy$ if and only if $x-y\in 2K$, i.e. if and only if the translates $x+K$ and $y+K$ intersect. The first and second steps in the above outline are accomplished via the following lemma.
Lemma 2.1 (cf. [CJMS23, Lemma 2.1]).
Suppose $d>10$ and $\Delta$ are so that $d^{12}<\Delta\leq\Delta_{K}$. Let $\Omega\subset\mathbb{R}^{d}$ be bounded and measurable. There exists some finite set of points $X\subset\Omega$ such that
  • $|X|\geq(1-2/d)(\Delta/2^{d})\operatorname{vol}(\Omega)$,
  • the graph $G(X,K)$ has maximum degree at most $\Delta+\Delta^{2/3}$, and
  • the graph $G(X,K)$ has maximum codegree at most $d^{-9}\Delta$.
(Here, the codegree of a pair of distinct vertices $(v_{1},v_{2})$ is the number of vertices $w$ of the graph for which both $v_{1}w$ and $v_{2}w$ are edges.) We prove Lemma 2.1 in Section 3, following the above outline. To accomplish the third step of the construction, we use the following graph-theoretic result.
Theorem 2.2 ([CJMS23, Theorem 1.3]).
Let $G$ be a graph on $n$ vertices with maximum degree at most $\Delta$, and suppose that the maximum codegree of any two vertices in $G$ is at most $\Delta/(2\log\Delta)^{7}$. Then there exists an independent set in $G$ of size at least
$(1-o(1))\frac{n\log\Delta}{\Delta},$
where the $o(1)$ tends to zero as $\Delta\to\infty$.
To complete the proof of Theorem 1.1, we must lower-bound $\Delta_{K}$; equivalently, we must give an upper bound on the volume of $I_{K}$. Indeed, the following holds.
Proposition 2.3.
There exists a constant $C_{\mathrm{int}}>0$ for which, for any centrally symmetric convex body $K\subset\mathbb{R}^{d}$ of volume $1$,
$\operatorname{vol}(I_{K})\leq\left(\frac{C_{\mathrm{int}}\log^{2}d}{d}\right)^{d/2}.$
It is in proving Proposition 2.3 that we will use the volume bounds of Schmuckenschläger and of Petty.
Proof of Theorem 1.1.
For each compact $\Omega\subset\mathbb{R}^{d}$, we find a finite set $A\subset\Omega$ so that
$|A|\geq(1-o(1))\frac{d\log d}{2^{d+1}}\operatorname{vol}(\Omega)$
and the balls $x+K$ and $y+K$ are disjoint for $x,y\in A$. Via a standard compactness argument (see [BMP05, Section 1.1]), having such a packing within each compact set is enough to give the desired bound on $\delta_{T}(K)$.
We may assume $d$ is sufficiently large: for $d$ small, we can use the trivial bound $\delta_{T}(K)\geq 2^{-d}$ and absorb the loss into the $o(1)$ term. Let $\Delta=\min\big{(}\Delta_{K},d^{d^{8/7}}\big{)}$, so that (using Proposition 2.3) we have $\Delta_{K}\geq\Delta>d^{12}$. So, by Lemma 2.1, we may choose a set $X\subset\Omega$ with
$|X|\geq\frac{(1-o(1))\Delta\operatorname{vol}(\Omega)}{2^{d}}$
so that $G(X,K)$ has maximum degree at most $\Delta+\Delta^{2/3}$ and maximum codegree at most $d^{-9}\Delta$. Since $\Delta\leq d^{d^{8/7}}$, we have $d^{-9}\Delta\leq\Delta/(2\log\Delta)^{7}$. So, we may apply Theorem 2.2 to the graph $G(X,K)$ to find an independent set $A\subset X$ with
$|A|\geq(1-o(1))\frac{|X|\log\Delta}{\Delta}\geq\frac{(1-o(1))\operatorname{vol}(\Omega)\log\Delta}{2^{d}}.$
This set $A$ is our set of centers. Finally, Proposition 2.3 implies that
$\log\Delta$ $\geq\min\left(d^{8/7}\log d,\log\Delta_{K}\right)$
$=\min\left(d^{8/7}\log d,\log\left((d\operatorname{vol}I_{K})^{-1}\right)\right)$
$\geq\frac{(1-o(1))d\log d}{2}.$
This finishes the proof of Theorem 1.1. ∎
2.3. Non-symmetric sets
Here, we explain how to deduce Corollary 1.2 from Theorem 1.1. This procedure goes back to Rogers and Shephard [RS57], who coupled their bound on the volume of $K-K$ with an observation of Minkowski [Min04] to attain the bound
$\delta_{T}(K)\geq(1-o(1))\frac{\sqrt{\pi d}}{2^{2d}}$
for any convex body $K$. For completeness, we outline the deduction.
Lemma 2.4 ([Min04]).
For any convex body $K\subset\mathbb{R}^{d}$,
$\frac{\delta_{T}(K)}{\operatorname{vol}K}=\frac{\delta_{T}\left(\frac{K-K}{2}\right)}{\operatorname{vol}\left(\frac{K-K}{2}\right)}.$
Proof sketch.
This equality follows from the fact that a discrete set $A\subset\mathbb{R}^{d}$ serves as the centers for a packing of $K$ if and only if it serves as the centers for a packing of $\frac{K-K}{2}$. ∎
Lemma 2.5 (Rogers–Shephard inequality; [RS57]).
For any convex body $K\subset\mathbb{R}^{d}$,
$\operatorname{vol}(K-K)\leq\binom{2d}{d}\operatorname{vol}K,$
with equality if and only if $K$ is a simplex.
We remark that, since equality in Lemma 2.5 holds in the case of a simplex $K=\triangle_{d}$, the trivial bound $\delta_{T}(\frac{\triangle_{d}-\triangle_{d}}{2})\leq 1$ furnishes the upper bound $\delta_{T}(\triangle_{d})\leq 2^{d}/\binom{2d}{d}\sim 2^{-d}\sqrt{\pi d}$ mentioned in the introduction.
Proof of Corollary 1.2.
Note that $L:=\frac{K-K}{2}$ is convex and centrally symmetric. So, chaining Lemma 2.4, Lemma 2.5, and Theorem 1.1 gives
$\delta_{T}(K)=\frac{\operatorname{vol}K}{\operatorname{vol}L}\delta_{T}(L)\leq\frac{2^{d}}{\binom{2d}{d}}\delta_{T}(L)\leq\frac{(1-o(1))d\log d}{2\binom{2d}{d}}.$
The corollary then follows from the asymptotic $\binom{2d}{d}=2^{2d}(\pi d)^{-1/2}(1+o(1))$. ∎
3. Selecting the set $X$: Proof of Lemma 2.1
We now prove Lemma 2.1. As previously mentioned, our strategy will be to first pick a set of points $\mathbf{X}\subset\Omega$ via a Poisson point process and then to remove a $o(1)$-fraction of the points to form $X$. The proof is quite similar to the proof of [CJMS23, Lemma 2.1]. The additional complexity is in bounding the maximum codegree, and this is where the choice of $\Delta_{K}$ as determined by $\operatorname{vol}(I_{K})$ is relevant. We will need the following standard tail bound for Poisson random variables.
Lemma 3.1 (Follows from [JŁR00, Theorem 2.1]).
A random variable $Z\sim\operatorname{Pois}(\lambda)$ satisfies the following upper tail bound: for each $t\geq 1$,
$\Pr\big{[}Z>(1+t)\lambda\big{]}\leq e^{-\lambda t/3}.$
We now show Lemma 2.1.
Proof of Lemma 2.1.
Through what follows, we write $I=I_{K}$, as the relevant body $K$ will not change. We choose the set $X$ in the following manner. Let $\mathbf{X}$ be a set of points chosen according to a Poisson point process inside $\Omega$ with intensity $\lambda:=2^{-d}\Delta$. Define the following subsets $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\subset\mathbf{X}$:
$\mathbf{X}_{1}$ =$=\left\{x\in\mathbf{X}:|\mathbf{X}\cap(x+2K)|>\Delta+\Delta^{2/3}\right\},$
$\mathbf{X}_{2}$ =$=\left\{x\in\mathbf{X}:\mathbf{X}\cap(x+2I)\text{ contains a point other than $x$}\right\},$
$S_{3}$ =$=\left\{(x,y)\in\mathbf{X}^{2}:x-y\not\in 2I\text{ and }|(\mathbf{X}\setminus\{x,y\})\cap(x+2K)\cap(y+2K)|\geq d^{-9}\Delta\right\},$
$\mathbf{X}_{3}$ =$=\left\{x\in\mathbf{X}:\exists y\in\mathbf{X}\text{ with }(x,y)\in S_{3}\right\}.$
Set $X=\mathbf{X}\setminus(\mathbf{X}_{1}\cup\mathbf{X}_{2}\cup\mathbf{X}_{3})$. We make the following observations.
  • By the definition of $\mathbf{X}_{1}$, the degree of any vertex of $G(X,K)$ is at most $\Delta+\Delta^{2/3}$.
  • By the definition of $\mathbf{X}_{2}$, each pair $(x,y)$ of distinct elements of $X$ satisfies $x-y\not\in 2I$.
  • By the definition of $\mathbf{X}_{3}$ and the above property, the maximum codegree of $G(X,K)$ is at most $d^{-9}\Delta$.
So, as long as $|X|\geq(1-2/d)(\Delta/2^{d})\operatorname{vol}(\Omega)$, $X$ will satisfy the conditions of the lemma. We will in fact show that $\mathbb{E}|X|$ exceeds this quantity; the result follows by Markov’s inequality. We have
$\mathbb{E}|\mathbf{X}|$ $=\lambda\operatorname{vol}(\Omega)=\Delta/2^{d}\cdot\operatorname{vol}(\Omega)\text{ and }$
$\mathbb{E}|X|$ $\geq\mathbb{E}|\mathbf{X}|-\mathbb{E}|\mathbf{X}_{1}|-\mathbb{E}|\mathbf{X}_{2}|-\mathbb{E}|S_{3}|.$
Hence, it suffices to show the following bounds:
(3.1) $\mathbb{E}|\mathbf{X}_{1}|$ $\leq e^{-d}\mathbb{E}|\mathbf{X}|$
(3.2) $\mathbb{E}|\mathbf{X}_{2}|$ $\leq d^{-1}\mathbb{E}|\mathbf{X}|$
(3.3) $\mathbb{E}|S_{3}|$ $\leq e^{-d}\mathbb{E}|\mathbf{X}|.$
Proof of (3.1).
Conditioned on $x\in\mathbf{X}$, $|\mathbf{X}\cap(x+2K))|-1$ is Poisson with mean
$\lambda\operatorname{vol}\big{(}(x+2K)\cap\Omega\big{)}\leq 2^{d}\lambda=\Delta.$
So, Lemma 3.1 implies
$\Pr[x\in\mathbf{X}_{1}|x\in\mathbf{X}]\leq e^{-(\Delta^{2/3}-1)/3}.$
This bound, together with the fact that $\Delta>d^{12}$, gives (3.1).
Proof of (3.2).
Conditioned on $x\in\mathbf{X}$, the random variable $|\mathbf{X}\cap(x+2I)|-1$ is Poisson with mean
$\lambda\operatorname{vol}\big{(}(x+2I)\cap\Omega\big{)}\leq\lambda\operatorname{vol}(2I)=\Delta\operatorname{vol}(2I)2^{-d}=\Delta\operatorname{vol}(I)\leq\frac{1}{d}$
So, $|\mathbf{X}\cap(x+2I)|>1$, i.e. $x\in\mathbf{X}_{2}$, with probability at most $1-e^{-1/d}\leq 1/d$. This gives (3.2).
Proof of (3.3).
If $(x,y)\in S_{3}$, then $x+2K$ and $y+2K$ must intersect, so $x-y\in 4K$. We now show that, for any pair of distinct points $(x,y)$ with $x-y\in 4K$,
$\Pr\left[(x,y)\in S_{3}\big{|}x,y\in\mathbf{X}\right]\leq e^{-d^{-10}\Delta}.$
Indeed, fixing $x$ and $y$ with $x-y\not\in 2I$ and conditioning on $x,y\in\mathbf{X}$, we have
$\big{|}(\mathbf{X}\setminus\{x,y\})\cap(x+2K)\cap(y+2K)\big{|}\sim\operatorname{Pois}\big{(}\lambda\operatorname{vol}((x+2K)\cap(y+2K))\big{)},$
and
$\operatorname{vol}\big{(}(x+2K)\cap(y+2K)\big{)}=2^{d}\operatorname{vol}\left(K\cap\left(K+\frac{x-y}{2}\right)\right)\leq 2^{d}d^{-10}$
since $(x-y)/2\not\in I$. We conclude that $|(\mathbf{X}\setminus\{x,y\})\cap(x+2K)\cap(y+2K)|$ is Poisson with mean at most $d^{-10}\Delta$. As a result, the probability that it exceeds $d^{-9}\Delta$ is at most, by Lemma 3.1,
$\Pr\left[\operatorname{Pois}(d^{-10}\Delta)\geq 4d^{-10}\Delta\right]\leq e^{-d^{-10}\Delta},$
as desired. Now, we can use this bound to upper-bound $\mathbb{E}|S_{3}|$:
$\mathbb{E}|S_{3}|$ $\leq e^{-d^{-10}\Delta}\cdot\mathbb{E}\big{|}\{(x,y):x,y\in\mathbf{X},\ x\neq y,\ x-y\in 4K\}\big{|}$
$\leq e^{-d^{-10}\Delta}\cdot\mathbb{E}|\mathbf{X}|\cdot\sup_{x\in\Omega}\mathbb{E}\left[\big{|}\{y\in\mathbf{X}\setminus\{x\}:x-y\in 4K\}\big{|}|x\in\mathbf{X}\right]$
$=e^{-d^{-10}\Delta}\cdot\mathbb{E}|\mathbf{X}|\cdot\lambda\operatorname{vol}(4K)=2^{d}\Delta e^{-d^{-10}\Delta}\cdot\mathbb{E}|\mathbf{X}|.$
The fact that $\Delta>d^{12}$ is enough to give (3.3).
4. Bounding the volume of $I_{K}$: Proof of Proposition 2.3
We now prove Proposition 2.3, i.e. that there exists a constant $C_{\mathrm{int}}>0$ so that, for every dimension $d$ and every centrally symmetric convex body $K$
$\operatorname{vol}(I_{K})\leq\left(\frac{C_{\mathrm{int}}\log^{2}d}{d}\right)^{d/2},$
where as before
$I_{K}=\left\{x\in\mathbb{R}^{d}:\operatorname{vol}(K\cap(K+x))>d^{-10}\right\}.$
To do this, we must define some auxiliary convex bodies. We will define these by their support functions: given a convex body $K\subset\mathbb{R}^{d}$, its support function is the function $h_{K}\colon\mathbb{R}^{d}\to\mathbb{R}$ defined by
$h_{K}(y)=\sup_{x\in K}x\cdot y.$
Definition 4.1.
Given a convex body $K\subset\mathbb{R}^{d}$, the polar body $K^{*}$ of $K$ is given by
$K^{*}:=\{y\in\mathbb{R}^{d}:|h_{K}(y)|\leq 1\}.$
The projection body $\Pi K$ of $K$ is defined so that, for each $x$ with $\|x\|_{2}=1$,
$h_{\Pi K}(x)=\operatorname{vol}_{\mathbb{R}^{d-1}}(\text{projection of $K$ onto }x^{\perp})$
The polar projection body, denoted $\Pi^{*}K$, is the polar of the projection body.
Our control on $I_{K}$ comes from the following result of Schmuckenschläger:
Lemma 4.2 ([Sch92, Theorem 1]).
For any symmetric convex body $K$ and any $\delta\in(0,1)$, one has
$(1-\delta)\Pi^{*}K\subset\{x\in\mathbb{R}^{d}:\operatorname{vol}(K\cap(K+x))>\delta\}\subset\log\left(\frac{1}{\delta}\right)\Pi^{*}K.$
We sketch a proof of the latter containment, which is the only part we need, using the theory of logarithmically concave (“log-concave”) distributions. (For the relevant definitions, see the introduction of [Pré73].) Our proof is essentially a streamlined version of Schmuckenschläger’s original proof.
Proof sketch.
Let $f(x)=\operatorname{vol}(K\cap(K+x))$. We have that $f$ is the probability density function of the convolution of the uniform measure on $K$ with itself. The uniform measure on $K$ is log-concave, as $K$ is convex; therefore, results of Prékopa [Pré73, Theorems 2 and 7] implies that $f$ is a log-concave function. Fix some vector $y\in\mathbb{R}^{d}$ with $\|y\|_{2}=1$, and define $g\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ by $g(t)=f(ty)$. Since $f$ is log-concave, $g$ is as well. Let $\pi$ be the projection operator onto $y^{\perp}$. We claim that $g^{\prime}(0)=-\operatorname{vol}(\pi(K))$. Indeed, this follows from the expression
[…]
The fact that $g$ is log-concave thus implies
$\log\operatorname{vol}(K\cap(K+ty))=\log g(t)\leq\log g(0)+t\frac{g^{\prime}(0)}{g(0)}=-t\cdot h_{\Pi K}(y)=-h_{\Pi K}(ty).$
Setting $x=ty$, we conclude that, if $\operatorname{vol}(K\cap(K+x))>\delta$, then $\log(1/\delta)\geq h_{\Pi K}(x)$. This is equivalent to $x\in(\log(1/\delta))\Pi^{*}K$. ∎
The second containment in Lemma 4.2 bounds $I_{K}$ by $\Pi^{*}K$ within a logarithmic factor in the dimension. It remains to upper-bound the volume of $\Pi^{*}K$. This can be done via an inequality of Petty.
Lemma 4.3 ([Pet71]).
Let $K\subset\mathbb{R}^{d}$ be a convex body, and let $B$ be a Euclidean ball satisfying $\operatorname{vol}K=\operatorname{vol}B$. Then
$\operatorname{vol}\Pi^{*}K\leq\operatorname{vol}\Pi^{*}B.$
In fact, much more can be said: the volume of $\Pi^{*}K$ is tightly controlled by the minimal surface area of a linear image of $K$; see [GP99, Theorem 3.6]. Petty’s inequality is enough for our purposes.
Proof of Proposition 2.3.
For each dimension $k\geq 1$, write $\gamma_{k}=\pi^{k/2}/(\frac{k}{2})!$ for the volume of the $k$-dimensional unit ball. [^3] For nonnegative real $x$, we use $x!$ as a shorthand for $\Gamma(x+1)$, where $\Gamma$ denotes the gamma function. Let $r_{k}$ be the radius of a unit-volume Euclidean ball in $\mathbb{R}^{k}$, so that $\gamma_{k}r_{k}^{k}=1$. Finally, let $B$ be the unit-volume Euclidean ball in $\mathbb{R}^{d}$. By Lemmas 4.2 and 4.3, we have
$\operatorname{vol}(I_{K})$ $=\operatorname{vol}\big{(}\{x\in\mathbb{R}^{d}:\operatorname{vol}(K\cap(K+x))>d^{-10}\big{)}$
$\leq(10\log d)^{d}\operatorname{vol}(\Pi^{*}K)$
(4.1) $\leq(10\log d)^{d}\operatorname{vol}(\Pi^{*}B).$
So, it remains to compute $\operatorname{vol}(\Pi^{*}B)$. For any $x\in\mathbb{R}^{d}$ of Euclidean norm $1$, the projection of $B$ onto $x^{\perp}$ is a $(d-1)$-dimensional ball of radius $r_{d}$, and so
$h_{\Pi B}(x)=\gamma_{d-1}r_{d}^{d-1}.$
As a result, $\Pi B$ is a Euclidean ball of radius $\gamma_{d-1}r_{d}^{d-1}$, and $\Pi^{*}B$ is a Euclidean ball of radius $(\gamma_{d-1}r_{d}^{d-1})^{-1}$. We conclude
(4.2) $\operatorname{vol}(\Pi^{*}B)=\frac{\gamma_{d}}{(\gamma_{d-1}r_{d}^{d-1})^{d}}=\frac{\gamma_{d}^{d}}{\gamma_{d-1}^{d}}=\left(\frac{\sqrt{\pi}\left(\frac{d-1}{2}\right)!}{\left(\frac{d}{2}\right)!}\right)^{d}\leq\left(\frac{2\pi}{d}\right)^{d/2},$
where we have used that $\gamma_{d-1}r_{d-1}^{d-1}=\gamma_{d}r_{d}^{d}=1$ and that
$\frac{x!}{\left(x-\frac{1}{2}\right)!}\geq\sqrt{x}$
for every real $x\geq 0$, which follows from the log-convexity of the gamma function and the property $x!/(x-1)!=x$. The result follows for $C_{\mathrm{int}}=200\pi$ by combining (4.1) and (4.2). ∎
Remark 4.4.
Instead of bounding $I_{K}$ inside a scaled image of $\Pi^{*}K$, one may obtain a cruder estimate by bounding $I_{K}$ inside a Euclidean ball, once a suitable linear transformation has been applied to $K$ (to obtain an isotropic body). [^4] As mentioned in the introduction, this was the argument given in an earlier version of the article. With this strategy, one may get a bound of the form
(4.3) $\operatorname{vol}(I_{K})^{1/d}\lesssim L_{K}\cdot\log d\cdot\operatorname{vol}\big{(}B_{\|\cdot\|_{2}}(0,1)\big{)}^{1/d}\lesssim\frac{\log d\cdot L_{K}}{\sqrt{d}},$
where $L_{K}$ is the slicing constant of $K$, the reciprocal of the largest volume of intersection of $K$ with any hyperplane. In our proof of Proposition 2.3, we were able to upper-bound all relevant quantities by the corresponding quantities for a Euclidean ball. However, $L_{K}$ is (nearly) minimized, rather than maximized, when $K$ is a Euclidean ball (see [KM22, pg. 203–206]). The question of giving constant upper bound on $L_{K}$ is the Bourgain slicing problem [Bou86, Bou87, Bal86], which was recently resolved by Klartag and Lehec [KL24], building on recent works [Che21, KL22, JLV22, Kla23, Gua24]. Combining this result with (4.3) is enough to show Proposition 2.3 and thus Theorem 1.1. (In fact, any bound of the form $L_{K}\lesssim\log^{O(1)}d$ is enough to prove Theorem 1.1, albeit with an $o(1)$ term which is larger by a constant factor.)
We note that the Bourgain slicing problem is closely related to other questions in convex geometry, such as the Kannan–Lovász–Simonovits conjecture [KLS95] and the thin-shell problem of Anttila, Ball, and Perissinaki [ABP03] . More examples of equivalent problems are listed in [MP89, Section 5], and more detail on the historical context of and progress on the slicing problem is given in the survey article [KM22].
Acknowledgements
The author would like to thank Jacob Fox for many helpful conversations and suggestions, Assaf Naor for communicating this proof of Proposition 2.3, and Dmitrii Zakharov for mentioning the connection between packing the simplex and the cross-polytope.